A population of iguanas on the Galapagos Islands has 20 individuals with webbed feet and 180 individuals with the dominant trait of non-webbed feet. What are the frequencies of webbed feet and non-webbed feet traits in this population?
Answer by checkley77(12844)

20/(20+180)=20/200=1/10 is the ratio of webbed to non-webbed feet iguanas.
180/200=9/10 is the ration of non-webbed to webbed feet iguanas.